Scores of an lQ test have a bell-shaped distribution with a mean of 100 and a standard deviation of 11. Use the empirical rule to determine the following. (a) What percentage of people has an IQ score between 89 and 1117 (b) What percentage of peopleasa Q score less than 78 or greater than 122? (C) What percentage of people has an IQ score greater than 111? (a) )%(Type an integer or a decimal.) (D)0% (Type an integer or a decimal ) (C) (Type an integer or a decimal.)
Answer #1

Given that, mean (mu) = 100 and standard deviation (sigma) = 11

According to empirical rule,

i) Approximately 68% of the data fall within 1 standard deviations of the mean.

ii) Approximately 95% of the data fall within 2 standard deviations of the mean.

iii) Approximately 99.7% of the data fall within 3 standard deviations of the mean.

ext { i) } mu -sigma = 100 - 11 = 89 ext { and } mu + sigma = 100+ 11 = 111

ii) μ-20 = 100 _ (2 * 11) 78 and μ + 2σ 100 (2 * 11) = 122

iii) μ-3σ-100一(3 * 11 )-67 and μ +3ơ 100 + (3 * 11) 1:33

Therefore,

a) 68% of the people has an IQ score between 89 and 111

b) 2.5% of the people has an IQ score less than 78 or greater than 122

c) 16% of the people has an IQ score greater than 111